On January 1, 2018, a machine was purchased for $120,000. The machine has an estimated salvage value of $9,600 and an estimated useful life of 5 years. The machine can operate for 120,000 hours before it needs to be replaced. The company closed its books on December 31 and operates the machine as follows: 2018, 24,000 hrs; 2019, 30,000 hrs; 2020, 18,000 hrs; 2021, 36,000 hrs; and 2022, 12,000 hrs.

Compute the annual depreciation charges over the machine’s life assuming a December 31 year-end for each of the following depreciation methods. (Round answers to 0 decimal places, e.g. 45,892.)

1. Straight-line Method $

Entry field with correct answer

2. Activity Method
Year
2018 $

Entry field with correct answer

2019 $

Entry field with correct answer

2020 $

Entry field with correct answer

2021 $

Entry field with correct answer

2022 $

Entry field with correct answer

3. Sum-of-the-Years'-Digits Method
Year
2018 $

Entry field with correct answer

2019 $

Entry field with correct answer

2020 $

Entry field with correct answer

2021 $

Entry field with correct answer

2022 $

Entry field with correct answer

4. Double-Declining-Balance Method
Year
2018 $

Entry field with correct answer

2019 $

Entry field with correct answer

2020 $

Entry field with correct answer

2021 $

Entry field with correct answer

2022 $

Entry field with incorrect answer

Assume a fiscal year-end of September 30. Compute the annual depreciation charges over the asset’s life applying each of the following methods. (Round answers to 0 decimal places, e.g. 45,892.)

Year

Straight-line Method

Sum-of-the-years'-digits method

Double-declining-balance method

2018 $

Entry field with correct answer

$

Entry field with correct answer

$

Entry field with incorrect answer

2019

Entry field with correct answer

Entry field with incorrect answer

Entry field with incorrect answer

2020

Entry field with correct answer

Entry field with incorrect answer

Entry field with incorrect answer now contains modified data

2021

Entry field with correct answer

Entry field with incorrect answer

Entry field with incorrect answer

2022

Entry field with correct answer

Entry field with incorrect answer now contains modified data

Entry field with incorrect answer

2023

Entry field with correct answer

Entry field with incorrect answer now contains modified data

Entry field with incorrect answer
Answer #1

1. straight line method = (cost – salvage value)/estimated life = (120000-9600)/5 = $22080

2. Activity method

2018 (120000-9600)/120000*24000

$22080

2019 (120000-9600)/120000*30000

$27600

2020 (120000-9600)/120000*18000

$16560

2021 (120000-9600)/120000*36000

$33120

2022 (120000-9600)/120000*12000

$11040

3. Sum-of-the-Years'-Digits Method

Sum-of-the-Years'-Digits = 5+4+3+2+1 = 15

2018 (120000-9600)*5/15

$36800

2019 (120000-9600)*4/15

$29440

2020 (120000-9600)*3/15

$22080

2021 (120000-9600)*2/15

$14720

2022 (120000-9600)*1/15

$7360

4. Double-Declining-Balance Method                     

Depreciation rate = 1/estimated life * 2 = 1/5 *2 = 40%

2018 (120000*40%)

$48000

2019 (120000-48000)*40%

$28800

2020 (120000-48000-28800)*40%

$17280

2021 (120000-48000-28800-17280)*40%

$10368

2022 ((120000-48000-28800-17280-10368)-9600

$5952

Part B

Straight-line Method

Sum-of-the-Years'-Digits Method

Double-Declining-Balance Method

2018

16560

27600

36000

2019

22080

31280

33600

2020

22080

23920

20160

2021

22080

16560

12096

2022

22080

9200

7258

2023

5520

1840

1286

22080*9/12 = 16560

22080*3/12 = 5520

(120000-9600)*5/15*9/12= 27600

((120000-9600)*5/15*3/12)+((120000-9600)*4/15*9/12) = 31280

((120000-9600)*4/15*3/12)+((120000-9600)*3/15*9/12) = 23920

((120000-9600)*3/15*3/12)+((120000-9600)*2/15*9/12) = 16560

((120000-9600)*2/15*3/12)+((120000-9600)*1/15*9/12) = 9200

((120000-9600)*1/15*3/12) = 1840

(120000*40%)*9/12 = 36000

(120000-36000)*40% = 33600

(120000-36000-33600)*40% = 20160

(120000-36000-33600-20160)*40% = 12096

(120000-36000-33600-20160-12096)*40% = 7258

(120000-36000-33600-20160-12096-7258)-9600 = 1286